OpenStudy (anonymous):

A yellow lamp emits light with a wavelength of 6.00 · 10-7 m. How many such photons are required to produce 10.0 joules worth of photons?

OpenStudy (shiraz14):

Use E=hc/(lambda) [where h is the Planck's constant (\[6.626X 10^{-34}\] J/s) and lambda is the wavelength in metres].

OpenStudy (shiraz14):

To answer your question, Energy of 1 photon = hc/(lambda) = (6.626X10^(-34))(3.0*10^8)/(6.00X10^(-7)) = 3.313X10^(-18) J Given that there are 10.0J of photons, Number of photons = 10.0/[3.313X10^(-18)] = 3.018X10^18 photons. This is assuming c = 3.0X10^8 m/s (although some texts might prefer 2.998X10^8 m/s).

OpenStudy (anonymous):

Thank you

OpenStudy (shiraz14):

You're welcome. And by the way, you might want to round the answer to the nearest integer (i.e. 3X10^18 photons) as some individuals might argue that there is no such thing as 0.018 photon (since a photon is currently widely presumed to be a discrete wave packet which cannot be divided any further).

OpenStudy (shiraz14):

Sorry, please ignore my previous reply - the 0.018 is in index form (I must have been dreaming when I typed this ... :p)